OpenStudy (anonymous):

Solve the following system of equations. Please show your work to receive full credit. x - y = 13 2x + y = 2

OpenStudy (ivettef365):

just add x - y = 13 2x + y = 2 ----------- ??

OpenStudy (anonymous):

x - y = 13 2x + y = 2 ( adding both the equations) ----------- 3x+0 =15 3x=15 x=15/3 x=5 Since x - y = 13 (substituting x=5 in this eq) 5- y = 13 5-13=y -8=y i.e. y=-8 thus {x,y}={5,-8}

OpenStudy (anonymous):

wow that was fast .. o.o